J8-16. A hydrogen-oxygen fuel cell delivers one ampere at 0.72 V with a fuel flow of 10.00 cm3/min. Independent measurement shows that 2.25 cm/min of fuel is escaping through the electrolyte unreacted. The cell operates in the standard state of 298°K and 1 atm. (a) Find the actual voltage and current efficiency. (b) Fid the magnitude of the three major contributions to heat transfer from the cell.
Answer #1

Efficiency=?v/v=(10-2.25)/10=0.775

Therefore actual voltage= 0.775*0.72=0.558v

Theagnitude requires potential data of other resources which are not given and therefore cannot be calculated.